OpenStudy (anonymous):

Simplify the expression \[\sqrt[4]{256a^2b^4}\]

OpenStudy (anonymous):

\[ \sqrt[4]{256a^2b^4}=\sqrt[4]{256} \cdot \sqrt[4]{a^2} \cdot \sqrt[4]{b^4} \]

OpenStudy (anonymous):

So what do I put as the answer? 4 square root a^2 b^4?

OpenStudy (anonymous):

do you know how to simplify it more?

OpenStudy (anonymous):

Nope I have no Idea how to do those

OpenStudy (anonymous):

do you know what \(\sqrt[4]{b^4}\) simplifies to?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

\[ \sqrt[4]{b^4} = |b| \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[ \sqrt[4]{a^2} = \sqrt{a} \]

OpenStudy (anonymous):

So the answer wil be 4 square root ab?

OpenStudy (anonymous):

\[ \sqrt[4]{256} = \sqrt[4]{2^8} = 2^2 = 4 \]

OpenStudy (anonymous):

Yeah basically.

OpenStudy (anonymous):

would you mind showing it to me in an equation?

OpenStudy (anonymous):

Unless you can assume \(b>0\) it's \(|b|\)